# Descent Methods for Vector Optimization Problems: A Majorization-Minimization Perspective

### v2

@@ 7 Second-order methods from a majorization-minimization perspective
$\alpha^{k}_{i}=\frac{\left\langle{\nabla F_{i}(x^{k})-\nabla F_{i}(x^{k-1}),x^{k}-x^{k-1}}\right\rangle}{\left\lVert{x^{k}-x^{k-1}}\right\rVert^{2}_{B_{k}}}$
for $i\in[m]$. Unfortunately, it is important to note that $\lambda^{k}$ is unavailable prior to computing the subproblem, and $\alpha^{k}$ is determined based on $B_{k}$. As an alternative, we can approximate $B_{k}$ using $\sum_{i\in[m]}\lambda^{k-1}_{i}{\nabla^{2}F_{i}(x^{k})}/{\alpha^{k-1}_{i}}$ instead of $\sum_{i\in[m]}\lambda^{k}_{i}{\nabla^{2}F_{i}(x^{k})}/{\alpha^{k}_{i}}$. Consequently, by updating $B_{k}$ via the BFGS method, the subproblem (27) coincides with that of preconditioned Barzilai-Borwein method for MOPs, which can be efficiently solved using the Frank-Wolfe method [9].
@@
Chen et al. [9] provided a detailed explanation of why the combination of the Barzilai-Borwein (BB) method and the BFGS method is essential in the multi-objective optimization context, particularly from a preconditioning perspective. Specifically, these two approaches function as complementary preconditioners for the search direction. While their effects tend to overlap in single-objective optimization, in the multi-objective setting, they fulfill distinct roles and demonstrate improved performance when applied together. In this section, we also highlight the indispensability of the Barzilai-Borwein rule in the quasi-Newton method from a majorization-minimization perspective.
@@ 8 Numerical Results
In this section, we present numerical results to demonstrate the performance of Barzilai-Borwein descent methods for VOPs (BBDVO) with polyhedral cones. We also compare BBDVO with steepest descent method for VOPs (SDVO) and equiangular direction method [28] for VOPs (EDVO). All numerical experiments were implemented in Python 3.7 and executed on a personal computer with an Intel Core i7-11390H, 3.40 GHz processor, and 16 GB of RAM. For all tested algorithms, we used Armijo line search (21) with $\sigma=10^{-4}$ and $\gamma=0.5$. To ensure that the algorithms terminate after a finite number of iterations, for all tested algorithms we used the stopping criterion $\|d(x)\|\leq 10^{-6}$. We also set the maximum number of iterations to 500. The test algorithms were executed on several test problems, and the problem illustration is given in Table 1. The dimensions of variables and objective functions are presented in the second and third columns, respectively. $x_{L}$ and $x_{U}$ represent lower bounds and upper bounds of variables, respectively. For each problem, we used the same initial points for different tested algorithms. The initial points were randomly selected within the specified lower and upper bounds. Dual subproblems of different algorithms were efficiently solved by Frank-Wolfe method. The recorded averages from the 200 runs include the number of iterations, the number of function evaluations, and the CPU time.
@@
For the tested problems, the partial order are induced by polyhedral cones +$\mathbb{R}^{2}_{+}$, $K_{1}$, and $K_{2}$, respectively, where
@@ 8.1 Numerical results for VOPs with $K=\mathbb{R}^{2}_{+}$
Table 2: Number of average iterations (iter), number of average function evaluations (feval), and average CPU time (time($ms$)) of SDVO, EDVO and BBDVO implemented on different test problems with $K=\mathbb{R}^{2}_{+}$
@@
BK1 1.00 2.00 0.21 1.00 2.00 0.39 27.87 32.11 2.37 1.00 1.00 0.23
DD1 71.09 199.63 12.14 43.14 163.92 10.77 354.26 354.26 33.48 7.49 8.91 1.26
Deb 57.86 376.40 13.95 55.14 350.18 16.50 10.18 26.02 1.42 4.41 6.58 0.84
FF1 28.95 29.11 3.97 16.00 61.75 4.40 9.07 10.00 0.93 4.91 6.13 0.78
Hil1 25.04 82.89 5.24 12.87 49.58 4.34 17.68 45.45 2.74 11.32 12.15 2.08
Imbalance1 88.23 178.31 11.86 387.91 388.32 40.39 405.13 405.92 34.90 2.61 3.54 0.46
JOS1a 311.23 311.23 34.89 16.53 18.05 3.47 65.67 65.67 7.87 1.00 1.00 0.29
LE1 22.44 52.57 3.40 7.98 29.78 1.70 11.34 29.26 1.35 4.55 7.03 0.81
PNR 11.33 48.80 2.13 12.23 15.97 1.76 13.65 21.89 1.41 4.18 4.74 0.66
WIT1 73.45 462.47 16.43 139.82 146.50 14.22 10.06 16.60 1.07 3.53 3.62 0.55
@@ 8.2 Numerical results for VOPs with $K=K_{1}$
[…]
Table 3: Number of average iterations (iter), number of average function evaluations (feval), and average CPU time (time($ms$)) of SDVO, EDVO and BBDVO implemented on different test problems with $K=K_{1}$
@@
BK1 1.00 2.65 0.23 7.49 9.85 1.18 43.66 47.15 4.04 1.00 1.00 0.16
DD1 93.30 376.98 18.65 118.15 120.10 14.63 495.62 495.62 48.61 42.07 47.34 6.61
Deb 102.47 878.36 29.21 27.88 146.52 6.47 68.61 145.74 8.80 35.69 71.88 6.55
FF1 33.20 105.30 6.25 17.24 56.85 3.45 40.00 57.24 4.35 16.09 17.03 2.51
Hil1 29.77 159.99 7.65 20.13 90.95 5.06 21.78 52.25 3.37 17.66 18.27 2.97
Imbalance1 77.34 309.32 13.53 397.37 397.44 44.40 492.52 492.52 42.95 28.78 31.18 4.49
JOS1a 81.13 81.13 9.69 6.57 17.06 2.53 106.12 106.12 12.63 1.00 1.00 0.24
LE1 14.02 69.28 3.01 12.37 55.82 2.45 15.63 39.36 1.89 6.29 7.47 1.09
PNR 30.34 178.44 6.61 13.10 38.79 2.38 19.95 25.23 2.05 9.57 10.77 1.65
WIT1 245.72 1095.05 71.75 198.45 245.46 22.39 405.71 408.04 38.08 151.35 156.72 21.66
@@ 8.3 Numerical results for VOPs with $K=K_{2}$
[…]
Table 4: Number of average iterations (iter), number of average function evaluations (feval), and average CPU time (time($ms$)) of SDVO, EDVO and BBDVO implemented on different test problems with $K=K_{2}$
@@
BK1 25.16 96.73 3.31 6.67 11.79 1.14 19.66 25.82 1.97 1.00 1.00 0.24
DD1 18.90 69.25 3.83 44.21 45.34 4.90 120.74 128.80 12.21 4.86 5.16 0.79
Deb 19.48 164.15 6.58 18.30 135.99 4.89 22.44 143.05 4.89 9.44 48.91 2.21
FF1 13.54 19.00 2.19 14.75 67.65 3.22 17.57 125.12 4.49 4.78 5.59 0.94
Hil1 15.40 79.01 4.08 17.99 84.51 4.54 15.74 39.73 2.47 8.24 9.09 1.51
Imbalance1 26.43 111.38 4.62 500.00 610.90 55.78 500.00 500.04 44.23 4.36 5.77 0.71
JOS1a 53.38 53.38 5.99 4.65 11.66 2.18 42.96 42.96 5.36 1.00 1.00 0.31
LE1 9.64 47.62 2.22 14.81 79.35 3.06 10.12 49.15 1.81 7.55 42.97 1.73
PNR 18.77 114.16 4.18 13.49 51.13 2.38 10.75 34.71 1.49 6.63 8.64 1.02
WIT1 24.11 147.56 5.89 308.99 973.42 44.31 10.07 28.11 1.42 8.42 9.98 1.34
@@
(a) BK1 (b) DD1 (c) FF1 (d) Imbalance1 (e) WIT1
Figure 1: Numerical results in value space obtained by BBDVO for problems BK1, DD1, FF1, Imbalance1 and WIT1 with $K=\mathbb{R}^{2}_{+}$, $K=K_{1}$ and $K=K_{2}$, respectively.
For test problems with different partial orders, the number of average iterations (iter), number of average function evaluations (feval), and average CPU time (time(ms)) of the different algorithms are listed in Tables 2, 3 and 4, respectively. We conclude that BBDVO outperforms SDVO and EDVO, especially for problems DD1 and Imbalance1. For SDVO, its performance is sensitive to the choice of transform matrix, changing the transform matrix in subproblem cannot improve the performance on all test problems. Naturally, a question arises that how to choose an appropriate transform matrix for a specific test problem in SDVO. It is worth noting that BBDVO can be viewed as SDVO with variable transform matrices ($\Lambda^{k}A$ is a transform matrix of $K$) and thus enjoys promising performance on these test problems. This provides a positive answer to the question. EDVO can also be viewed as SDVO with variable transform matrices, it generates descent directions with norm less than $1$ (the minimizer of subproblem is the minimal norm element of the convex hull of some unit vectors), decelerating the convergence in large-scale problems (the initial point may be far from the Pareto set).
Fig. 1 plots the final points obtained by BBDVO on problems BK1, DD1, FF1, Imbalance1 and WIT1 with $K=\mathbb{R}^{2}_{+}$, $K=K_{1}$ and $K=K_{2}$, respectively. We can observe that enlarging the partial order cone reduces the number of obtained Pareto critical points, especially in the long tail regions, where improving one objective function slightly can sacrifice the others greatly. As a result, we can use an order cone containing the non-negative orthant in real-world MOPs to obtain Pareto points with a better trade-off.

### v3

@@ 6.3 Backtracking method for VOPs with polyhedral cones
$\left\lVert{x^{k+1}-x^{*}}\right\rVert\leq\sqrt{1-\gamma\min\limits_{i\neq j,i,j\in[l]}\left\{\frac{\left\langle{A_{i},\bm{\mu}}\right\rangle}{\left\langle{A_{i},\bm{\ell}}\right\rangle}\frac{\left\langle{A_{j},\bm{\mu}}\right\rangle}{\left\langle{A_{j},\bm{\ell}}\right\rangle}\right\}}\left\lVert{x^{k}-x^{*}}\right\rVert$
holds for all $k\geq 0$.
Proof.
By Lemma 6.2(i), we have $A\bm{\mu}\preceq\alpha^{k}\preceq A\bm{\ell}$. For any $i\neq j$, consider the choice $\alpha^{k}_{i}=\left\langle{A_{i},\bm{\mu}}\right\rangle$ and $\alpha^{k}_{j}=\left\langle{A_{j},\bm{\ell}}\right\rangle$, then we have
$\min\limits_{i\in[l]}\frac{\alpha^{k}_{i}}{\left\langle{A_{i},\bm{\ell}}\right\rangle}\min\limits_{i\in[l]}\frac{\left\langle{A_{i},\bm{\mu}}\right\rangle}{\alpha^{k}_{i}}\leq\frac{\left\langle{A_{i},\bm{\mu}}\right\rangle}{\left\langle{A_{i},\bm{\ell}}\right\rangle}\frac{\left\langle{A_{j},\bm{\mu}}\right\rangle}{\left\langle{A_{j},\bm{\ell}}\right\rangle}.$
The arbitrary of $i$ and $j$ with $i\neq j$ yields
$\min\limits_{i\in[l]}\frac{\alpha^{k}_{i}}{\left\langle{A_{i},\bm{\ell}}\right\rangle}\min\limits_{i\in[l]}\frac{\left\langle{A_{i},\bm{\mu}}\right\rangle}{\alpha^{k}_{i}}\leq\min\limits_{i\neq j,i,j\in[l]}\left\{\frac{\left\langle{A_{i},\bm{\mu}}\right\rangle}{\left\langle{A_{i},\bm{\ell}}\right\rangle}\frac{\left\langle{A_{j},\bm{\mu}}\right\rangle}{\left\langle{A_{j},\bm{\ell}}\right\rangle}\right\}.$
Conversely, notice that
$\min\limits_{i\in[l]}\frac{\alpha^{k}_{i}}{\left\langle{A_{i},\bm{\ell}}\right\rangle}\geq\min\limits_{i\in[l]}\frac{\left\langle{A_{i},\bm{\mu}}\right\rangle}{\left\langle{A_{i},\bm{\ell}}\right\rangle}$   and    $\min\limits_{i\in[l]}\frac{\left\langle{A_{i},\bm{\mu}}\right\rangle}{\alpha^{k}_{i}}\geq\min\limits_{i\in[l]}\frac{\left\langle{A_{i},\bm{\mu}}\right\rangle}{\left\langle{A_{i},\bm{\ell}}\right\rangle}$,
and no single $\alpha^{k}$ can simultaneously satisfy both equilities unless there exists two distinct indices $i_{0}\neq j_{0}$ such that
$\min\limits_{i\in[l]}\frac{\left\langle{A_{i},\bm{\mu}}\right\rangle}{\left\langle{A_{i},\bm{\ell}}\right\rangle}=\frac{\left\langle{A_{i_{0}},\bm{\mu}}\right\rangle}{\left\langle{A_{i_{0}},\bm{\ell}}\right\rangle}=\frac{\left\langle{A_{j_{0}},\bm{\mu}}\right\rangle}{\left\langle{A_{j_{0}},\bm{\ell}}\right\rangle},$
which yields
$\min\limits_{i\in[l]}\frac{\alpha^{k}_{i}}{\left\langle{A_{i},\bm{\ell}}\right\rangle}\min\limits_{i\in[l]}\frac{\left\langle{A_{i},\bm{\mu}}\right\rangle}{\alpha^{k}_{i}}\geq\min\limits_{i\neq j,i,j\in[l]}\left\{\frac{\left\langle{A_{i},\bm{\mu}}\right\rangle}{\left\langle{A_{i},\bm{\ell}}\right\rangle}\frac{\left\langle{A_{j},\bm{\mu}}\right\rangle}{\left\langle{A_{j},\bm{\ell}}\right\rangle}\right\}.$
Hence,
$\min\limits_{i\in[l]}\frac{\alpha^{k}_{i}}{\left\langle{A_{i},\bm{\ell}}\right\rangle}\min\limits_{i\in[l]}\frac{\left\langle{A_{i},\bm{\mu}}\right\rangle}{\alpha^{k}_{i}}=\min\limits_{i\neq j,i,j\in[l]}\left\{\frac{\left\langle{A_{i},\bm{\mu}}\right\rangle}{\left\langle{A_{i},\bm{\ell}}\right\rangle}\frac{\left\langle{A_{j},\bm{\mu}}\right\rangle}{\left\langle{A_{j},\bm{\ell}}\right\rangle}\right\}.$
The desired result follows by Lemma 6.2(ii) and (iv). ∎
In contrast to the line search method for VOPs, the corresponding method for SOPs preserves the same linear convergence rate as its counterpart without line search. This discrepancy arises because the line search in VOPs is imposed with respect to a strict dominance relation by the underlying partial order. We clarify the distinction with the following example.
Example 6.1.
Let $K=\mathbb{R}^{2}_{+}$ and $A=I_{2}$. Consider first the case where $\alpha^{k}=(\bm{\mu}_{1},\bm{\ell}_{2})^{T}$. As illustrated in Fig. 2(a), the ratio $\alpha^{k}/t_{k}$ lies on the blue line determined by the strict dominance relation. In this situation, the line search results in a slower linear convergence rate, with the worst-case rate being $\mathcal{O}\left(\left(\sqrt{1-\gamma\frac{\bm{\mu}_{1}\bm{\mu}_{2}}{\bm{\ell}_{1}\bm{\ell}_{2}}}\right)^{k}\right)$.
(a) Line search for VOPs
(b) Line search for SOPs
Figure 2: Differences in line search for VOPs and SOPs.
In contrast, if $\alpha^{k}$ falls on the red point in Fig. 2(a), then $\alpha^{k}/t_{k}$ lies on the red line. In this case, line search for VOPs potentially improves the practical linear convergence rate, with the worst-case rate coinciding with that of the method without line search. As illustrated in Fig. 2(b), line search for SOPs tends to improve linear convergence rate, and the worst case rate is the same as that of its counterpart without line search.
[…]
Data: $x^{0}\in\mathbb{R}^{n},~A\in\mathcal{A},~0\prec\bm{\ell}^{0}\preceq A\bm{\ell},~\tau>1$
1 for $k=0,1,$... do
2   Update $\alpha^{k}:=\bm{\ell}^{k}$
3   Update $x^{k+1}:=\mathop{\arg\min}\limits_{x\in\mathbb{R}^{n}}\max\limits_{\lambda\in\Delta^{\alpha^{k}}_{l}}\left\{\left\langle{\lambda,AJF(x^{k})(x-x^{k})}\right\rangle+\frac{1}{2}\left\lVert{x-x^{k}}\right\rVert^{2}\right\}$
4 if $x^{k+1}=x^{k}$ then
5    return $K$-stationary point $x^{k}$
6 else
7    $s_{i}=0,i\in[l]$
8    repeat
9         Update $\alpha_{i}^{k}=\tau^{s_{i}}\bm{\ell}^{k}_{i},i\in[l]$
10         Update $x^{k+1}:=\mathop{\arg\min}\limits_{x\in\mathbb{R}^{n}}\max\limits_{\lambda\in\Delta^{\alpha^{k}}_{l}}\left\{\left\langle{\lambda,AJF(x^{k})(x-x^{k})}\right\rangle+\frac{1}{2}\left\lVert{x-x^{k}}\right\rVert^{2}\right\}$
11       for $i=1,\cdots,l$ do
12          if $\left\langle{A_{i},F(x^{k+1})-F(x^{k})}\right\rangle>\left\langle A_{i},JF(x^{k})(x^{k+1}-x^{k})\right\rangle+\frac{\alpha^{k}_{i}}{2}\|x^{k+1}-x^{k}\|^{2}$ then
13              Update $s_{i}=s_{i}+1$
14            end if
15
16         end for
17
18    until $\left\langle{A_{i},F(x^{k+1})-F(x^{k})}\right\rangle\leq\left\langle A_{i},JF(x^{k})(x^{k+1}-x^{k})\right\rangle+\frac{\alpha^{k}_{i}}{2}\|x^{k+1}-x^{k}\|^{2},~i\in[l]$;
19   end if
20  Update $\bm{\ell}^{k+1}:=\alpha^{k}/\tau$
21 end for
22
Algorithm 9 Backtracking method for VOPs with polyhedral cones
We consider the following surrogate:
$G_{k,A,\alpha^{k}}(x):=AJF(x^{k})(x-x^{k})+\frac{1}{2}\left\lVert{x-x^{k}}\right\rVert^{2}\alpha^{k}.$ (30)
We can show that $G_{k,A,\alpha^{k}}(\cdot)$ is a non-majorizing surrogate function of $A(F(\cdot)-F(x^{k}))$ near $x^{k}$.
Proposition 6.5.
Let $G_{k,A,\alpha^{k}}(\cdot)$ be defined as (30). Then, the following statements hold.
  • (i)
    $A(F(x^{k+1})-F(x^{k}))\preceq G_{k,A,\alpha^{k}}(x^{k+1})$.
  • (ii)
    If $F(\cdot)$ is $K$-convex, then $G_{k,A,\alpha^{k}}(\cdot)\in\mathcal{S}_{\alpha^{k},\alpha^{k}}(AF,x^{k})$.
  • (iii)
    If $F(\cdot)$ is strongly $K$-convex with $\bm{\mu}\in{\rm int}(K)$, then $G_{k,A,\alpha^{k}}(\cdot)\in\mathcal{S}_{\alpha^{k}-A\bm{\mu},\alpha^{k}}(AF,x^{k})$.
Proof.
Lemma 6.3.
  • (i)
    $\alpha^{k}\prec\tau A\bm{\ell}$;
  • (ii)
    $\{x^{k}\}$ converges to an efficient solution $x^{*}$ of (VOP);
  • (iii)
    $\left\lVert{x^{k+1}-x^{*}}\right\rVert\leq\sqrt{1-\tau\min_{i\in[l]}{\left\langle{A_{i},\bm{\mu}}\right\rangle}/{\left\langle{A_{i},\bm{\ell}}\right\rangle}}\left\lVert{x^{k}-x^{*}}\right\rVert,~\forall k\geq 0$.
Proof.
(i) Suppose, to the contrary, that $\alpha^{k}_{i}\geq\tau\left\langle{A_{i},\bm{\ell}}\right\rangle$ holds for some $i\in[l]$. Then the backtracking procedure would be triggered only when $\alpha^{k}_{i}\geq\left\langle{A_{i},\bm{\ell}}\right\rangle$, which contradicts the backtracking condition. Assertions (ii) and (iii) can be obtained by using the similar arguments as in the proof of Lemma 4.1. ∎
In contrast to the line search method for VOPs, the backtracking method preserves the same linear convergence rate as its counterpart without line search. This discrepancy arises because backtracking for VOPs is imposed with respect to a weak dominance relation induced by the underlying partial order. We clarify the distinction with the following example.
Example 6.2.
Let $K=\mathbb{R}^{2}_{+}$ and $A=I_{2}$. Consider the case where $\bm{\ell}^{k}=(\bm{\mu}_{1},\bm{\ell}_{2})^{T}$.
Figure 3: Differences between line search and backtracking for VOPs.
As illustrated in Fig. 3, the ratio $\bm{\ell}^{k}/t_{k}$ lies on the blue line determined by the strict dominance relation. In this situation, the line search yields a slower linear convergence rate, with the worst-case rate given by $\mathcal{O}\left(\left(\sqrt{1-\gamma\frac{\bm{\mu}_{1}\bm{\mu}_{2}}{\bm{\ell}_{1}\bm{\ell}_{2}}}\right)^{k}\right)$. In contrast, the $\alpha^{k}$ lies on the red line determined by the weak dominance relation. In this case, the backtracking procedure for VOPs preserves the same linear convergence rate as its counterpart without line search.
@@ 20  Update $\bm{\ell}^{k+1}:=\alpha^{k}/\tau$
A notable advantage of backtracking is its ability to adapt to the local smoothness and to preserve the same linear convergence rate as its counterpart without line search. However, this benefit comes at the price of increased per-iteration computational cost, since backtracking requires repeatedly solving the associated subproblems.
@@ 7 Numerical Results
In this section, we present numerical results to demonstrate the performance of Barzilai-Borwein descent methods for VOPs (BBDVO) with polyhedral cones. We also compare BBDVO with steepest descent method for VOPs (SDVO) and equiangular direction method [28] for VOPs (EDVO). All numerical experiments were implemented in Python 3.7 and executed on a personal computer with an Intel Core i7-11390H, 3.40 GHz processor, and 16 GB of RAM.
7.1 Implementation details
For all tested algorithms, we used Armijo line search (27) with $\sigma=10^{-4}$ and $\gamma=0.5$. The test algorithms were executed on several test problems, and the problem illustration is given in Table 1. The dimensions of variables and objective functions are presented in the second and third columns, respectively. $x_{L}$ and $x_{U}$ represent lower bounds and upper bounds of variables, respectively.
@@ 7.1 Implementation details
For the tested problems, the partial order are induced by polyhedral cones $\mathbb{R}^{2}_{+}$, $K_{1}$, and $K_{2}$, respectively, where
@@
The polyhedral cones are illustrated in Fig. 4.
Figure 4: Illustration of the polyhedral cones.
For each problem, we used the same initial points for different tested algorithms. The initial points were randomly selected within the specified lower and upper bounds. Dual subproblems of different algorithms were efficiently solved by Frank-Wolfe method. To guarantee a fair comparison, we decided to let the algorithms run until one of the following stopping conditions was satisfied:
  • 1.
    the current solution satisfies $\left\lVert{\bar{d}^{k}}\right\rVert\leq 10^{-6}$, where $\bar{d}$ is the steepest descent direction (20) with $A=\bar{A}$ and $\bar{A}\in\mathcal{A}$ with row vectors $\left\lVert{\bar{A}_{i}}\right\rVert=1$ for i=1,,li=1,...,l;
  • 2.
    the number of iterations reaches 500.
The recorded averages from the 200 runs include the number of iterations, the number of function evaluations, and the CPU time. The performance profiles [12] in terms of iterations, function evaluations and CPU time are used to illustrate the overall performance of the 200 runs.
@@ 7.2 Numerical results for VOPs with $K=\mathbb{R}^{2}_{+}$
Table 2: Number of average iterations (iter), number of average function evaluations (feval), and average CPU time (time(msms)) of SDVO, EDVO and BBDVO implemented on different test problems with $K=\mathbb{R}^{2}_{+}$
@@
BK1 1.00 2.00 0.24 36.93 40.15 4.03 30.12 34.07 2.67 1.00 1.00 0.21
DD1 70.95 199.40 12.14 248.33 250.81 30.41 376.96 376.96 38.11 7.33 8.51 1.24
Deb 57.79 376.02 13.67 5.67 13.15 1.36 10.48 26.39 1.43 4.51 6.67 0.79
FF1 28.95 29.11 3.93 8.22 9.15 1.38 8.68 9.51 1.03 4.68 5.90 0.84
Hil1 24.89 82.53 5.15 15.62 36.52 3.05 20.09 51.60 3.06 11.42 12.25 2.07
Imbalance1 88.23 178.31 12.35 387.91 388.32 43.47 420.04 420.99 42.01 2.62 3.60 0.50
JOS1a 302.72 302.72 35.80 16.51 18.03 3.40 74.13 74.44 9.29 1.00 1.00 0.26
LE1 22.44 52.57 3.72 7.98 29.78 1.70 11.80 29.93 1.47 4.65 7.13 0.82
PNR 11.24 48.40 2.83 12.32 16.09 1.85 14.37 23.24 1.75 4.28 4.77 0.71
WIT1 73.30 461.37 19.93 139.88 146.58 15.15 10.64 17.97 1.33 3.59 3.68 0.62
(a) Iterations
(b) Function Evaluations
(c) CPU Time
Figure 5: Performance profiles on the test problems in Table 2 with $K=\mathbb{R}^{2}_{+}$.
@@ 7.3 Numerical results for VOPs with $K=K_{1}$
In this case, we denote the set of transform matrices $\mathcal{A}_{1}:=\{A:AK_{1}=\mathbb{R}^{2}_{+}\}$. For SDVO, we choose $A^{1}=\begin{pmatrix}5&\ -1\\ -1&\ 5\end{pmatrix}\in\mathcal{A}_{0}$ and $\hat{A}^{1}\in\mathcal{A}_{1}$ in subproblem, respectively, where $\hat{A}^{1}:=\{A:A_{i}=A^{1}_{i}/\max\{1,\left\lVert{\nabla F_{i}(x^{0})}\right\rVert_{\infty}\},~i=1,2\}$.
Table 3: Number of average iterations (iter), number of average function evaluations (feval), and average CPU time (time(msms)) of SDVO, EDVO and BBDVO implemented on different test problems with $K=K_{1}$
@@
BK1 1.00 2.65 0.21 7.51 9.87 1.21 48.30 51.56 4.07 1.00 1.00 0.22
DD1 92.40 373.32 18.42 118.35 120.30 15.34 496.99 496.99 53.13 41.85 47.12 6.39
Deb 102.21 876.19 27.68 27.82 146.28 6.29 81.45 161.00 9.44 35.69 71.88 6.48
FF1 33.19 105.29 5.85 17.23 56.82 3.16 45.35 65.10 4.73 16.00 16.95 2.61
Hil1 28.28 151.71 7.15 19.41 86.98 4.71 23.09 56.07 3.51 17.74 18.35 3.10
Imbalance1 77.34 309.32 13.55 397.37 397.44 46.97 500.00 500.00 50.14 28.78 31.18 4.49
JOS1a 69.86 69.86 6.82 6.57 17.06 2.55 117.30 117.43 13.78 1.00 1.00 0.26
LE1 14.02 69.28 2.88 12.37 55.82 2.46 15.99 39.60 1.92 6.31 7.49 1.14
PNR 27.10 157.72 5.71 12.76 37.65 2.15 21.80 27.54 2.41 9.78 10.97 1.64
WIT1 244.44 2084.16 111.19 199.03 245.92 24.42 414.29 416.78 42.52 158.78 164.91 22.80
(a) Iterations
(b) Function Evaluations
(c) CPU Time
Figure 6: Performance profiles on the test problems in Table 2 with $K=K_{1}$.
@@ 7.4 Numerical results for VOPs with $K=K_{2}$
We denote the set of transform matrices $\mathcal{A}_{2}:=\{A:AK_{2}=\mathbb{R}^{2}_{+}\}$. For SDVO, we choose $A^{2}=\begin{pmatrix}5&\ 1\\ 1&\ 5\end{pmatrix}\in\mathcal{A}_{0}$ and $\hat{A}^{2}\in\mathcal{A}_{2}$ in subproblem, respectively, where $\hat{A}^{2}:=\{A:A_{i}=A^{2}_{i}/\max\{1,\left\lVert{\nabla F_{i}(x^{0})}\right\rVert_{\infty}\},~i=1,2\}$.
Table 4: Number of average iterations (iter), number of average function evaluations (feval), and average CPU time (time(msms)) of SDVO, EDVO and BBDVO implemented on different test problems with $K=K_{2}$
@@
BK1 22.29 85.27 2.52 6.67 11.79 1.25 20.69 26.53 1.92 1.00 1.00 0.21
DD1 17.16 62.14 3.37 46.38 47.71 5.03 133.89 134.00 13.33 4.84 5.29 0.83
Deb 19.48 164.15 5.43 18.22 135.69 4.63 21.95 141.27 4.40 9.47 48.94 2.02
FF1 13.52 18.80 1.96 14.76 67.80 3.04 17.68 122.48 4.04 4.72 5.74 0.91
Hil1 14.69 75.19 3.86 17.60 82.61 4.44 16.61 42.27 2.65 8.38 9.24 1.52
Imbalance1 25.76 105.22 4.60 500.00 610.90 61.64 500.00 500.04 48.91 4.35 5.76 0.73
JOS1a 46.11 46.11 4.47 4.58 10.16 1.77 48.35 51.67 6.34 1.00 1.00 0.26
LE1 9.64 47.62 2.05 14.79 79.00 3.08 11.40 54.09 1.95 7.58 43.00 1.63
PNR 17.22 103.38 3.60 13.30 50.20 2.41 11.12 36.16 1.69 6.80 8.83 1.04
WIT1 23.12 140.31 5.32 311.78 975.91 44.38 10.46 29.67 1.52 8.66 9.95 1.27
(a) Iterations
(b) Function Evaluations
(c) CPU Time
Figure 7: Performance profiles on the test problems in Table 2 with $K=K_{2}$.
@@
(a) BK1 (b) FF1 (c) Hil1 (d) PNR (e) WIT1
Figure 8: Numerical results in value space obtained by BBDVO for problems BK1, FF1, Hil1, PNR and WIT1 with $\mathbb{R}^{2}_{+}, K=K$$K_{1}$ and $K=K_{2}$, respectively.
For test problems with different partial orders, the number of average iterations (iter), number of average function evaluations (feval), and average CPU time (time(ms)) of the different algorithms are listed in Tables 2, 3 and 4, respectively. We conclude that BBDVO outperforms SDVO and EDVO, especially for problems DD1 and Imbalance1. For SDVO, its performance is sensitive to the choice of transform matrix, changing the transform matrix in subproblem cannot improve the performance on all test problems. Naturally, a question arises that how to choose an appropriate transform matrix for a specific test problem in SDVO. It is worth noting that BBDVO can be viewed as SDVO with variable transform matrices ($\Lambda^{k}A$ is a transform matrix of $K$) and thus enjoys promising performance on these test problems. This provides a positive answer to the question. EDVO can also be viewed as SDVO with variable transform matrices, it generates descent directions with norm less than $1$ (the minimizer of subproblem is the minimal norm element of the convex hull of some unit vectors), decelerating the convergence in large-scale problems (the initial point may be far from the Pareto set). Figs. 5, 6 and 7 present the performance profiles based on iterations, function evaluations, and CPU time. The results confirm that the proposed BBDVO significantly outperforms SDVO and EDVO.
Fig. 8 plots the final points obtained by BBDVO on problems BK1, FF1, Hil1, PNR and WIT1 with $\mathbb{R}^{2}_{+}, K=K$$K_{1}$ and $K=K_{2}$, respectively. We can observe that enlarging the partial order cone reduces the number of obtained Pareto critical points, especially in the long tail regions, where improving one objective function slightly can sacrifice the others greatly. As a result, we can use an order cone containing the non-negative orthant in real-world MOPs to obtain Pareto points with a better trade-off.